OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1)

OpenStudy (anonymous):

the one that looks like \[y-y_1=m(x-x_1)\] if \(m=4,x_1=-3,y_1=-1\)

OpenStudy (anonymous):

wait what

OpenStudy (anonymous):

what "what"? make the replacements with the numbers

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\]\[y-(-1)=4(x-(-3))\]\[x+1=4(x+3)\]

OpenStudy (anonymous):

oops i meant \[y+1=4(x+3)\] with a \(y\)

OpenStudy (anonymous):

haha thank you!